OpenStudy (wendy.ivette11714):

A 15ft ladder is used to clean out the gutters on a house. The angle the ladder forms with the ground is 60 degrees. How far up the house is the ladder? A. 13 B.15 C.17 Will medal!!

OpenStudy (wendy.ivette11714):

@KendrickLamar2014 @just_one_last_goodbye @Elsa213

OpenStudy (wendy.ivette11714):

Kim is stalking me ;)

OpenStudy (wendy.ivette11714):

Why did you delete that? e.e

OpenStudy (hotpinkgirl):

idk

OpenStudy (wendy.ivette11714):

I'm getting no help ;-;

OpenStudy (kendricklamar2014):

Let me see..

OpenStudy (wendy.ivette11714):

I think I'm gonna close it

OpenStudy (kendricklamar2014):

I havnt dont these types of problems in a long time :(

OpenStudy (kendricklamar2014):

@johnweldon1993 can you help her?

OpenStudy (johnweldon1993):

Sure! Draw a diagram of the situation  We have a right triangle with the angle of 60 known...we have the hypotenuse as 15...and we want to know the opposite side of the triangle...what trig function uses opposite and hypotenuse?

OpenStudy (johnweldon1993):

*or let me know if you haven't used trig before...there are other methods as well :)

OpenStudy (wendy.ivette11714):

Sine

OpenStudy (johnweldon1993):

Okay good, trig it is And right \(\large sin(\theta) = \frac{opposite}{hypotenuse}\) So we know \(\large \theta = 60\) and we know we want to solve for the opposite...so we just use algebra to write this as \[\large opposite = 15 * sin(60)\] Which comes out to what?

OpenStudy (kendricklamar2014):

So you would multiply 15 * sin(60)? Is that correct? @johnweldon1993

OpenStudy (johnweldon1993):

Correct @KendrickLamar2014

OpenStudy (wendy.ivette11714):

-4.57?

OpenStudy (johnweldon1993):

Make sure your calculator is in "degrees" mode...not radians :)

OpenStudy (wendy.ivette11714):

Is it B?

OpenStudy (johnweldon1993):

It is not...dont guess though it's very easy Hint* plugging sin(60) into your calculator in degrees mode...comes out to a result of roughly 0.866

OpenStudy (kendricklamar2014):

Multiply: \[15 \times 0.866\]

OpenStudy (wendy.ivette11714):

A.

OpenStudy (johnweldon1993):

There ya go!

OpenStudy (wendy.ivette11714):

Thank you!!

OpenStudy (johnweldon1993):

Not a problem!

OpenStudy (kendricklamar2014):

Thanks for refreshing my memory on this @johnweldon1993 . I appreciate your help :)